OpenStudy (haileyd):

x+7y=-14 -2x+14y=0 Solve by elimination.

OpenStudy (chosenmatt):

needhelp?

OpenStudy (anonymous):

the key to elimination is that you want to be able to cancel out one variable from each of the equations. x+7y=-14 -2x+14y=0 none of the terms added together will cancel out... for example: x + (-2x)= -x ****you want this to equal 0 or 7y+14y= 21y ******does not equal to 0 so we must multiply one equation so that we can do that! let's multiply the first equation by 2 and see what can cancel out... 2(x+7y) = 2(-14) ~~~ 2x+14y= -28 now we have these 2 equations 2x+14y=-28 -2x+14y=0 now add the equations together we get: 28y= -28 (the x will cancel out and you end up with what y will equal!) y=-1 now we plug this -1 into any of the 2 equations in order to get our x value :)

OpenStudy (haileyd):

thank you, that helped me some^_^